W={(x1, x2, x3)∈R3: x2+x3=0}. Find two subspace W1,W2 of R such that R3= W⊕W1 and R3=W⊕W2 but W1≠W2
1
Expert's answer
2020-05-25T21:11:00-0400

Let denote "V=\mathbb{R}^3" be a real vector space.

Given that "W=\{(x_1,x_2,x_3) \in V|\: x_2+x_3=0\}" subspace of "V" .

Let, "W_1=\{(x_1,x_1,x_1)\in V : x_1\in\mathbb{R}\}" and "W_2=\{(2x_1,-x_1,x_1)\in V: \: x_1\in \mathbb{R}\}" ,clearly, "W,W_1,W_2" all different as a set.

Claim 1: "W_1 ,W_2" are subspace of "V"

Proof:

Let, "a,b,c,d \in \mathbb{R}" and for every "u,v \in W_1 \: \&\: w,y \in W_2" consider the linear combination as follows,

Case-I: clearly, "u=(u_1,u_1,u_1)\&v=(v_1,v_1,v_1)" such that,


"au+bv=(au_1+bv_1,au_1+bv_1,au_1+bv_1)"


which implies,

"au+bv\in W_1"

Hence "W_1" is vector subspace.


Case-II: In a similar manner as case-I , consider the

"cw+dy\in W_2"

Thus our claimed is proved.


Claim 2: "V=W \oplus W_1"

Proof:

Clearly, we see that

"V=W+ W_1"

Thus we will show that "W\cap W_1=\{0\}" . Suppose for any "w\in W \cap W_1" such that "w\neq0" ,thus "w\in W \& w\in W_1" which implies,

"(x_1,x_2,-x_2)=(x_1,x_1,x_1)\\\implies (x_1,x_2,x_3)=(0,0,0)"

Hence we arrived at contradiction that "w\neq 0" .

Thus,

"V=W\oplus W_1"

We are done.


Claim 3: "V=W\oplus W_2"

Proof:

Observe that "W+W_2=\{(3x_1,x_2-x_1,x_3+x_1): x_1,x_2,x_3 \in \mathbb{R}\}\supset \mathbb{R}^3" which implies


"V=W+W_2 \hspace{1cm}(\because W+W_1\subset V)"

Now, we will show that

"W\cap W_2=\{0\}"

Clearly, exactly applying the same arguments as in claim 2 we get,"W\cap W_2=\{0\}" .

Thus,

"V=W\oplus W_2"

Hence, we are done.